There is no need to develop a Simio model as part of this homework. You can certainly use a Simio model to validate the results that you obtain using the spreadsheet. For an M/M/1 queue with mean interarrival time 1.25 minutes and mean service time 1 minute, find all five of Wa, W, La, L, and p. For each, interpret in words. Be sure to state all your units (always!), and the relevant time frame of operation. 1. 2. Based on the arrival and service time data in problem 1, suppose that wed like to see what would happen if the arrival rate were to increase in small steps; maybe a single server barbershop would like to increase business by some advertising or coupons. Create a spreadsheet or a computer program, and re-evaluate all five of Wa, W, La, L, and p, except increasing the arrival rate by 5% over its original value, then by 15% over its original value, then by 20% over its original value, and so on up through increasing it by 100% over its original value (i.e., doubling the arrival rate). Make plots of each of the five metrics as functions of the percent increase in the arrival rate. Discuss. Repeat Problem 1 for an M/M/3 queue with mean interarrival time 1.25 minutes and mean service time 3 minutes at each of the three servers. Hint: You might want to consider creating a computer program or spreadsheet for this problem 3. 4. In Problem 3, increase the arrival rate in the same steps as in Problem 2, and again evaluate all ofWo, W, Lo,L, and p at each step. If you run into any particular problems at some point as you increase the arrival rate, how many more servers (beyond the three that you already have) would you need to calm things down?
Answer #1

1)

l =1.25 mins or 48 customers per hour

u =1 min or 60 customers per hour

Wq =Ws- u = l/(u(u-l)) . = 48/60(12) = 1/15 hours = 4 mins

Ws =Ls/l = =1/(u-l) = 1/(12) =5 mins

Lq =l Wq = l^2/ u(u-l)) = 48*48/ 60 *12 = 16/5 = 3.2

Ls =l/u-l= 48/ (60-48) = 4 patients

rho = l/u =48/60= .8

b)

in minutes in rate terms 5% inc 10% inc 15% inc 20% inc 50%inc 100%inc
l 1.25 48 50.4 52.8 55.2 57.6 72 96
u 1 60 60 60 60 60 60 60
base 5% inc 10% inc 15% inc 20% inc 50%inc 100%inc
wq (mins) 4 5.25 7.333333333 11.5 24 -6 -2.666666667
ws (mins) 5 6.25 8.333333333 12.5 25 -5 -1.666666667
ls 3.2 4.41 6.453333333 10.58 23.04 -7.2 -4.266666667
lq 4 5.25 7.333333333 11.5 24 -6 -2.666666667
rho 0.8 0.84 0.88 0.92 0.96 1.2 1.6

3)

1)

l =1.25 mins or 48 customers per hour

u =3 min or 20 customers per hour per server

c =3 servers

po= 1/(1+rho rho^2/2 + rho^c/ c! (1- rho/c)) = 1/(1+.8 + .8^/2 +.8^3/ 6(1- .8/3)) =.058

Wq = lq/l = .0539

Ws = wq +1/u =

Lq = rho^(c+1)/(c-1)! (c-rho)^2) Po = 2.5888

L = lWq + l/u =

rho =l/u =48/20 =2.4

utilization= l/cu =48/3*20= .8

b)

in minutes in rate terms 5% inc 10% inc 15% inc 20% inc 50%inc 100%inc
l 1.25 48 50.4 52.8 55.2 57.6 72 96
u 3 20 20 20 20 20 20 20
3 base 5% inc 10% inc 15% inc 20% inc 50%inc 100%inc
wq (mins) 0.053932584 0.07431789 0.108608731 0.177637726 0.385570844 -0.116546763 -0.063681592
ws (mins) 0.103932584 0.12431789 0.158608731 0.227637726 0.435570844 -0.066546763 -0.013681592
ls 4.988764045 6.265621656 8.374541004 12.56560249 25.08888062 -4.791366906 -1.313432836
lq 2.588764045 3.745621656 5.734541004 9.805602492 22.20888062 -8.391366906 -6.113432836
rho 2.4 2.52 2.64 2.76 2.88 3.6 4.8
p0 0.056179775 0.042799058 0.030599755 0.019466615 0.009297136 -0.035971223 -0.074626866
utilization 0.8 0.84 0.88 0.92 0.96 1.2 1.6

please use the formulas mention in the text I am unable to add spreadsheet to this